# Two-Timescale Gradient Descent Ascent Algorithms for Nonconvex Minimax Optimization

### v1

@@ 2 Related Work
The investigation of nonconvex-nonconcave minimax optimization problems has become a central topic in machine learning community, inspired by the advent of generative adversarial networks (GANs) (Goodfellow et al., 2014) and adversarial learning (Madry et al., 2018; Sinha et al., 2018). In this context, most works focus on defining a notion of goodness or the development of new procedures for reducing oscillations (Daskalakis and Panageas, 2018; Adolphs et al., 2019; Jin et al., 2020) and speeding up the convergence of gradient dynamics (Heusel et al., 2017; Balduzzi et al., 2018; Mertikopoulos et al., 2019). For example, Daskalakis and Panageas (2018) showed that the stable fixed points of GDA are not necessarily saddle points, while Adolphs et al. (2019) proposed Hessian-based algorithms where their stable fixed points are saddle points. In this context, Balduzzi et al. (2018) also developed a symplectic gradient adjustment algorithm for finding stable fixed points in potential games and Hamiltonian games. Notably, Heusel et al. (2017) have proposed TTGDA and TTSGDA formally for training GANs and showed that the saddle points are also stable fixed points of a continuous limit of these algorithms under strong conditions. All of these convergence results are either local or asymptotic and thus can not cover our nonasymptotic results for TTGDA and TTSGDA in the nonconvex-concave settings. Mertikopoulos et al. (2019) has provided nonasymptotic guarantees for a special class of nonconvex-nonconcave minimax optimization problems under variational stability and/or the so-called Minty condition. However, while these conditions must hold in convex-concave setting, they do not necessarily hold in the nonconvex-concave setting. Subsequent to our ICML work (Lin et al., 2020a), we are aware of great effort that has been devoted to understanding the nonconvex-nonconcave minimax optimization (Yang et al., 2020a; Diakonikolas et al., 2021; Liu et al., 2021; Lee and Kim, 2021; Ostrovskii et al., 2021a; Fiez et al., 2022; Jiang and Chen, 2023; Grimmer et al., 2023; Daskalakis et al., 2023; Zheng et al., 2023; Cai et al., 2024). Some of their results are valid for the nonconvex-concave setting but can not cover our results in this paper. This is an active research area and we refer the interested reader to the aforementioned articles for more details.
@@ 3 Preliminaries and Technical Background
We present the basic definitions for Lipschitz functions and smooth functions as follows,
@@
There are another notion of stationarity based on $\nabla f$ in smooth and nonconvex-concave setting (Nouiehed et al., 2019; Lu et al., 2020; Lin et al., 2020b; Xu et al., 2023). A pair of points $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m}\times\mathcal{Y}$ is an $\epsilon$-stationary point of $f$ if $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|\leq\epsilon$ and $\|\mathbf{y}^{+}-\mathbf{y}\|\leq\frac{\epsilon}{\ell}$ where $\mathbf{y}^{+}$ is defined by
@@
Intuitively, the term $\ell\|\mathbf{y}^{+}-\mathbf{y}\|$ stands for the distance between $\mathbf{y}\in\mathcal{Y}$ and $\mathbf{y}^{+}$ obtained by performing a one-step projected partial gradient ascent at $(\mathbf{x},\mathbf{y})$ starting from $\mathbf{y}$. This is also akin to the norm of gradient mapping (Nesterov, 2013). As highlighted by Thekumparampil et al. (2019), this notion is weaker than Definitions 3.3 and 3.8. For the sake of completeness, we provide a complete characterization of the relationship between these notions by showing that they can be translated in both directions with the extra computational cost. To pursue this goal, we introduce a new notion of stationarity based on $\nabla f$ as follows,
@@
The existing stationarity of $f$ is equivalent to Definition 3.10 up to a constant. Indeed, if $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m}\times\mathcal{Y}$ is an $\epsilon$-stationary point as per Definition 3.10, we have $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|\leq\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y}^{+})\|+\ell\|\mathbf{y}^{+}-\mathbf{y}\|\leq 2\epsilon$. Conversely, if $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m}\times\mathcal{Y}$ satisfies $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|\leq\epsilon$ and $\|\mathbf{y}^{+}-\mathbf{y}\|\leq\frac{\epsilon}{\ell}$, we have $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y}^{+})\|\leq\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|+\ell\|\mathbf{y}^{+}-\mathbf{y}\|\leq 2\epsilon$.
We focus on the notion as per Definition 3.10 and relate it to the stationarity of $\Phi$ (cf. Definition 3.3 and 3.8). The following propositions summarize the details.
@@
Propositions 3.11 and 3.12 show that one needs to pay additional $O(\kappa\log(1/\epsilon))$ or $O(\epsilon^{-2})$ gradient evaluations for translating the stationarity of $f$ to that of $\Phi$ in nonconvex-strongly-concave and nonconvex-concave settings. In this sense, the notions of stationarity as per Definitions 3.3 and 3.8 are stronger than the notion of stationarity as per Definition 3.10.
@@
Subsequent to our ICML work, Ostrovskii et al. (2021b) identified an error in the proof of Lin et al. (2020a, Proposition 4.12) and proposed a new notion of stationarity based on $\nabla f$ which is stronger than the existing stationarity of $f$. After communicating with the authors of Ostrovskii et al. (2021b), we agree that Lin et al. (2020a, Proposition 4.12) (i.e., the precursor of Proposition 3.12) is still valid and the error can be fixed by using the notion in Definition 3.10. We also acknowledge that their notion is strictly stronger than ours in Definition 3.10 when $\mathcal{Y}\neq\mathbb{R}^{n}$ while being tractable from a computational viewpoint.
@@ 4 Smooth Minimax Optimization
We prove the bounds on the complexity of solving smooth and nonconvex-concave minimax optimization problems. Our focuses are TTGDA and TTSGDA (see Algorithms 1 and 2) with a proper choice of $(\eta_{\mathbf{x}}^{t},\eta_{\mathbf{y}}^{t})_{t\geq 0}$; recall that these methods were shown to converge locally and yield empirical success in practice (Heusel et al., 2017). Intuitively, the reason for this phenomenon is that the choice of $\eta_{\mathbf{x}}^{t}\ll\eta_{\mathbf{y}}^{t}$ matches the asymmetric structure of nonconvex-concave minimax optimization problems.
@@ 4.1 Main results
In addition to smoothness condition, the function $f(\mathbf{x},\mathbf{y})$ is assumed to be concave in $\mathbf{y}$ but nonconvex in $\mathbf{x}$. Imposing such one-side concavity is necessary since Daskalakis et al. (2021) has shown that it is NP-hard to determine if an approximate min-max point exists in general smooth and nonconvex-nonconcave setting. Imposing the boundedness condition for $\mathcal{Y}$ is for practical purpose since the ambiguity sets are generally bounded in real application problems (Sinha et al., 2018). Relaxing this condition requires more complicated arguments and we leave it to future work.
@@
Theorem 4.6 is the analog of Theorem 4.4 in the nonconvex-concave setting but with two different results. First, the order of the ratio becomes $\eta_{\mathbf{y}}/\eta_{\mathbf{x}}=\Theta(\epsilon^{4})$ which highlights highly asymmetric structure here. Second, the theoretical results are valid for the case of $M=1$ and in that case they are different from that derived in Theorem 4.4.
@@ 4.2 Discussions
Our complexity results do not contradict the classical results on the divergence of GDA with fixed stepsize in the convex-concave setting. Indeed, there are a few key distinctions: (1) our results guarantee that the proposed algorithm visits an $\epsilon$-stationary point at some iterate, not the last iterate; (2) our results guarantee the stationarity of some iterate $\mathbf{x}_{t}$ rather than $(\mathbf{x}_{t},\mathbf{y}_{t})$. As such, our proof permits the possibility of significant changes in $\mathbf{y}_{t}$ even when $\mathbf{x}_{t}$ is very close to stationarity. This, together with our choice $\eta_{\mathbf{x}}^{t}\ll\eta_{\mathbf{y}}^{t}$, removes the contradiction. Based on the above fact, we highlight that our algorithms can be used to obtain an approximate saddle point in smooth and convex-concave setting (i.e., optimality for both $\mathbf{x}$ and $\mathbf{y}$). Instead of averaging, we can run two passes of TTGDA for minimax and maximin optimization problems separately. We use $\eta_{\mathbf{x}}^{t}\ll\eta_{\mathbf{y}}^{t}$ in the first pass while we use $\eta_{\mathbf{x}}^{t}\gg\eta_{\mathbf{y}}^{t}$ in the second pass. Either pass returns an approximate stationary point for each variable $\mathbf{x}$ and/or $\mathbf{y}$, which is defined using the corresponding Moreau envelop and which jointly forms an approximate saddle point.
@@ 4.3 Proof sketch
$\delta_{t}\leq(1-\tfrac{1}{2\kappa}+4\kappa^{3}\ell^{2}(\eta_{\mathbf{x}}^{t})^{2})\delta_{t-1}+4\kappa^{3}(\eta_{\mathbf{x}}^{t})^{2}\|\nabla\Phi(\mathbf{x}_{t-1})\|^{2}.$
@@
By viewing TTGDA as an inexact gradient descent for minimizing the function $\Phi(\cdot)$, we derive the key descent inequality as follows,
@@
We have that $\mathcal{Y}^{\star}(\mathbf{x})$ will not be a singleton for each $\mathbf{x}$ and $\mathcal{Y}^{\star}(\mathbf{x}_{1})$ can be dramatically different from $\mathcal{Y}^{\star}(\mathbf{x}_{2})$ even when $\mathbf{x}_{1},\mathbf{x}_{2}$ are close to each other. This implies that $\min_{\mathbf{y}\in\mathcal{Y}^{\star}(\mathbf{x}_{t})}\|\mathbf{y}_{t}-\mathbf{y}\|$ is no longer a viable error to control.
@@ 5 Nonsmooth Minimax Optimization
We prove the bounds on the complexity of solving nonsmooth and nonconvex-concave minimax optimization problems. Indeed, we assume that $f(\mathbf{x},\mathbf{y})$ is $L$-Lipschitz and $\rho$-weakly convex in $\mathbf{x}$. Our focuses are also TTGDA and TTSGDA with a proper choice of $(\eta_{\mathbf{x}}^{t},\eta_{\mathbf{y}}^{t})_{t\geq 0}$.
@@ 5.1 Main results
The settings defined by Assumptions 5.1 and 5.2 generalizes the nonsmooth optimization and are intuitively more challenging than smooth minimax optimization studied in Section 4. The following lemma is an analog of Lemma 4.3.
@@
Theorem 5.4 shows that the adaptive choice of $\eta_{\mathbf{y}}^{t}$ better exploits the nonconvex-strongly-concave structure in the nonsmooth setting; indeed, $\eta_{\mathbf{y}}^{t}=1/(\mu t)$ for $t\in\{1,2,\ldots,B\}$ which corresponds to the first epoch. Then, for each of subsequent epochs, we set the same sequence of stepsizes. Such finding is consistent with that for (stochastic) subgradient-type methods for minimizing a strongly convex function (Nemirovski and Yudin, 1983; Rakhlin et al., 2012; Shamir and Zhang, 2013). We remark that the complexity bounds for TTGDA and TTSGDA are the same in terms of $1/\epsilon$, which is also consistent with the classical results for nonsmooth optimization (Nemirovski and Yudin, 1983; Davis and Drusvyatskiy, 2019).
@@
Theorems 5.4 and 5.6 provide the gradient complexity results for TTGDA and TTSGDA when applied to solve the nonsmooth and nonconvex-(strongly)-concave minimax optimization problems. The orders of $O(\epsilon^{-4})$ and $\Theta(\epsilon^{-4})$ reflect the effect of nonsmoothness when compared to that presented in Theorem 4.4 and 4.6. Furthermore, we guarantee that an $\epsilon$-stationary point is visited within a certain number of iterations rather than guaranteeing last-iterate convergence. Finally, Theorem 5.4 is valid for the case of $M=1$ and such result is different from the theoretical results in Theorem 4.4.

### v2

@@ 2 Related Work
The investigation of nonconvex-nonconcave minimax optimization problems has become a central topic in the machine learning community, inspired by the advent of generative adversarial networks (GANs) (Goodfellow et al., 2014) and adversarial learning (Madry et al., 2018; Sinha et al., 2018). In this context, most works focus on defining a notion of goodness or the development of new procedures for reducing oscillations (Daskalakis and Panageas, 2018; Adolphs et al., 2019; Jin et al., 2020) and speeding up the convergence of gradient dynamics (Heusel et al., 2017; Balduzzi et al., 2018; Mertikopoulos et al., 2019). For example, Daskalakis and Panageas (2018) showed that the stable fixed points of GDA are not necessarily saddle points, while Adolphs et al. (2019) proposed Hessian-based algorithms where their stable fixed points are saddle points. In this context, Balduzzi et al. (2018) also developed a symplectic gradient adjustment algorithm for finding stable fixed points in potential games and Hamiltonian games. Notably, Heusel et al. (2017) have proposed TTGDA and TTSGDA formally for training GANs and showed that the saddle points are also stable fixed points of a continuous limit of these algorithms under strong conditions. All of these convergence results are either local or asymptotic and thus can not cover our nonasymptotic results for TTGDA and TTSGDA in the nonconvex-concave settings. Mertikopoulos et al. (2019) has provided nonasymptotic guarantees for a special class of nonconvex-nonconcave minimax optimization problems under variational stability and/or the so-called Minty condition. However, while these conditions must hold in convex-concave setting, they do not necessarily hold in the nonconvex-concave setting. Subsequent to our earlier work (Lin et al., 2020a), we are aware of significant effort that has been devoted to understanding the nonconvex-nonconcave minimax optimization problem (Yang et al., 2020a; Diakonikolas et al., 2021; Liu et al., 2021; Lee and Kim, 2021; Ostrovskii et al., 2021a; Fiez et al., 2022; Jiang and Chen, 2023; Grimmer et al., 2023; Daskalakis et al., 2023; Zheng et al., 2023; Cai et al., 2024). Some of their results are valid for the nonconvex-concave setting but can not cover our results in this paper. This is an active research area and we refer the interested reader to the aforementioned articles for more details.
@@ 3 Preliminaries and Technical Background
We present standard definitions for Lipschitz functions and smooth functions.
@@
There is another notion of stationarity based on $\nabla f$ in smooth and nonconvex-concave setting (Nouiehed et al., 2019; Lu et al., 2020; Lin et al., 2020b; Xu et al., 2023). A pair of points $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m}\times\mathcal{Y}$ is an $\epsilon$-stationary point of $f$ if $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|\leq\epsilon$ and $\|\mathbf{y}^{+}-\mathbf{y}\|\leq\frac{\epsilon}{\ell}$ where $\mathbf{y}^{+}$ is defined by
@@
Intuitively, the term $\ell\|\mathbf{y}^{+}-\mathbf{y}\|$ stands for the distance between $\mathbf{y}\in\mathcal{Y}$ and $\mathbf{y}^{+}$ obtained by performing a one-step projected partial gradient ascent at $(\mathbf{x},\mathbf{y})$ starting from $\mathbf{y}$. This is akin to the norm of the gradient mapping (Nesterov, 2013). As highlighted by Thekumparampil et al. (2019), this notion is weaker than Definitions 3.3 and 3.8. For the sake of completeness, we provide a complete characterization of the relationship between these notions by showing that they can be translated in both directions with the extra computational cost. To pursue this goal, we introduce a new notion of stationarity based on $\nabla f$ as follows.
@@
The stationarity of $f$ is equivalent to Definition 3.10 up to a constant. Indeed, if $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m}\times\mathcal{Y}$ is an $\epsilon$-stationary point as per Definition 3.10, we have $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|\leq\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y}^{+})\|+\ell\|\mathbf{y}^{+}-\mathbf{y}\|\leq 2\epsilon$. Conversely, if $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m}\times\mathcal{Y}$ satisfies $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|\leq\epsilon$ and $\|\mathbf{y}^{+}-\mathbf{y}\|\leq\frac{\epsilon}{\ell}$, we have $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y}^{+})\|\leq\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|+\ell\|\mathbf{y}^{+}-\mathbf{y}\|\leq 2\epsilon$.
We focus on the notion in Definition 3.10 and relate it to the stationarity of $\Phi$ (cf. Definition 3.3 and 3.8). The following propositions summarize the details.
@@
Propositions 3.11 and 3.12 show that one needs to pay an additional $O(\kappa\log(1/\epsilon))$ or $O(\epsilon^{-2})$ gradient evaluations for translating the stationarity of $f$ to that of $\Phi$ in nonconvex-strongly-concave and nonconvex-concave settings. In this sense, the notions of stationarity as per Definitions 3.3 and 3.8 are stronger than the notion of stationarity as per Definition 3.10.
@@
Ostrovskii et al. (2021b) identified an error in the proof of Lin et al. (2020a, Proposition 4.12) and proposed a new notion of stationarity based on $\nabla f$ which is stronger than the existing stationarity of $f$. After communicating with the authors of Ostrovskii et al. (2021b), we agree that Lin et al. (2020a, Proposition 4.12) (i.e., the precursor of Proposition 3.12) is still valid and the error can be fixed by using the notion in Definition 3.10. We also acknowledge that their notion is strictly stronger than ours in Definition 3.10 when $\mathcal{Y}\neq\mathbb{R}^{n}$ while being tractable from a computational viewpoint.
@@ 4 Smooth Minimax Optimization
We prove the bounds on the complexity of solving smooth and nonconvex-concave minimax optimization problems. Our focus is TTGDA and TTSGDA (see Algorithms 1 and 2) with a proper choice of $(\eta_{\mathbf{x}}^{t},\eta_{\mathbf{y}}^{t})_{t\geq 0}$; recall that these methods were shown to converge locally and are useful in practice (Heusel et al., 2017). Intuitively, the reason for this phenomenon is that the choice of $\eta_{\mathbf{x}}^{t}\ll\eta_{\mathbf{y}}^{t}$ matches the asymmetric structure of nonconvex-concave minimax optimization problems.
@@ 4.1 Main results
In addition to smoothness condition, the function $f(\mathbf{x},\mathbf{y})$ is assumed to be concave in $\mathbf{y}$ but nonconvex in $\mathbf{x}$. Imposing such one-side concavity is necessary since Daskalakis et al. (2021) have shown that it is NP-hard to determine if an approximate min-max point exists in a general smooth and nonconvex-nonconcave setting. Imposing the boundedness condition for $\mathcal{Y}$ is an assumption of convenience that reflects the fact that the ambiguity sets are generally bounded in real applications (Sinha et al., 2018). Relaxing this condition requires more complicated arguments and we leave it to future work.
@@
Theorem 4.6 is the analog of Theorem 4.4 in the nonconvex-concave setting but with two different results. First, the order of the ratio becomes $\eta_{\mathbf{y}}/\eta_{\mathbf{x}}=\Theta(\epsilon^{4})$ which highlights the highly asymmetric structure here. Second, the theoretical results are valid for the case of $M=1$ and in that case they are different from that derived in Theorem 4.4.
@@ 4.2 Discussions
Our complexity results do not contradict the classical results on the divergence of GDA with fixed stepsize in the convex-concave setting. Indeed, there are a few key distinctions: (1) our results guarantee that the proposed algorithm visits an $\epsilon$-stationary point at some iterate, not the last iterate; (2) our results guarantee the stationarity of some iterate $\mathbf{x}_{t}$ rather than $(\mathbf{x}_{t},\mathbf{y}_{t})$. As such, our proof permits the possibility of significant changes in $\mathbf{y}_{t}$ even when $\mathbf{x}_{t}$ is very close to stationarity. This, together with our choice $\eta_{\mathbf{x}}^{t}\ll\eta_{\mathbf{y}}^{t}$, removes the contradiction. Based on the above fact, we highlight that our algorithms can be used to obtain an approximate saddle point in the smooth and convex-concave setting (i.e., optimality for both $\mathbf{x}$ and $\mathbf{y}$). Instead of averaging, we can run two passes of TTGDA for minimax and maximin optimization problems separately. We use $\eta_{\mathbf{x}}^{t}\ll\eta_{\mathbf{y}}^{t}$ in the first pass while we use $\eta_{\mathbf{x}}^{t}\gg\eta_{\mathbf{y}}^{t}$ in the second pass. Either pass returns an approximate stationary point for each variable $\mathbf{x}$ and/or $\mathbf{y}$, which is defined using the corresponding Moreau envelope and which jointly forms an approximate saddle point.
@@ 4.3 Proof sketch
$\delta_{t}\leq(1-\tfrac{1}{2\kappa}+4\kappa^{3}\ell^{2}(\eta_{\mathbf{x}}^{t})^{2})\delta_{t-1}+4\kappa^{3}(\eta_{\mathbf{x}}^{t})^{2}\|\nabla\Phi(\mathbf{x}_{t-1})\|^{2},$
@@
By viewing TTGDA as inexact gradient descent for minimizing the function $\Phi(\cdot)$, we derive the key descent inequality as follows,
@@
We note that $\mathcal{Y}^{\star}(\mathbf{x})$ will not be a singleton for each $\mathbf{x}$ and $\mathcal{Y}^{\star}(\mathbf{x}_{1})$ can be dramatically different from $\mathcal{Y}^{\star}(\mathbf{x}_{2})$ even when $\mathbf{x}_{1},\mathbf{x}_{2}$ are close to each other. This implies that $\min_{\mathbf{y}\in\mathcal{Y}^{\star}(\mathbf{x}_{t})}\|\mathbf{y}_{t}-\mathbf{y}\|$ is no longer a viable error to control.
@@ 5 Nonsmooth Minimax Optimization
We prove the bounds on the complexity of solving nonsmooth and nonconvex-concave minimax optimization problems. Indeed, we assume that $f(\mathbf{x},\mathbf{y})$ is $L$-Lipschitz and $\rho$-weakly convex in $\mathbf{x}$. Our focus is TTGDA and TTSGDA with a proper choice of $(\eta_{\mathbf{x}}^{t},\eta_{\mathbf{y}}^{t})_{t\geq 0}$.
@@ 5.1 Main results
The setting defined by Assumptions 5.1 and 5.2 generalizes the nonsmooth optimization and are intuitively more challenging than smooth minimax optimization studied in Section 4. The following lemma is an analog of Lemma 4.3.
@@
Theorem 5.4 shows that the adaptive choice of $\eta_{\mathbf{y}}^{t}$ better exploits the nonconvex-strongly-concave structure in the nonsmooth setting; indeed, $\eta_{\mathbf{y}}^{t}=1/(\mu t)$ for $t\in\{1,2,\ldots,B\}$ which corresponds to the first epoch. Then, for each subsequent epoch, we set the same sequence of stepsizes. Such a finding is consistent with that for (stochastic) subgradient-type methods for minimizing a strongly convex function (Nemirovski and Yudin, 1983; Rakhlin et al., 2012; Shamir and Zhang, 2013). We remark that the complexity bounds for TTGDA and TTSGDA are the same in terms of $1/\epsilon$, which is also consistent with the classical results for nonsmooth optimization (Nemirovski and Yudin, 1983; Davis and Drusvyatskiy, 2019).
@@
Theorems 5.4 and 5.6 provide the gradient complexity results for TTGDA and TTSGDA when applied to solve the nonsmooth and nonconvex-(strongly)-concave minimax optimization problems. The orders of $O(\epsilon^{-4})$ and $\Theta(\epsilon^{-4})$ reflect the effect of nonsmoothness when compared to that presented in Theorem 4.4 and 4.6. Furthermore, we guarantee that an $\epsilon$-stationary point is visited within a certain number of iterations rather than guaranteeing last-iterate convergence. Finally, Theorem 5.4 is valid for the case of $M=1$ and such a result is different from the theoretical results in Theorem 4.4.